OpenStudy (anonymous):

Is ( x- 3) a factor of x^3 + 4x^2 + 2?

OpenStudy (mathmath333):

if \(\Large ( x- 3)\) is a factor of \(\Large\rm \begin{align} \color{black}{x^3 + 4x^2 + 2}\end{align}\) then it must satisfy \(\Large \begin{align} \color{black}{f(3)=x^3 + 4x^2 + 2=0}\end{align}\)